# Axion-like Dark Matter Search with Space-based Gravitational Wave Detectors

### v1

@@ IV.3 Expected sensitivity
where $P_{\mathrm{oms}}$ and $P_{\mathrm{acc}}$ are the noise amplitude parameters. For the BBO, the noises PSDs are given by Corbin and Cornish (2006)
$S_{\mathrm{oms}}^{\mathrm{BBO}}(f)$ $=\frac{2.0\times 10^{-34}}{3L^{2}}\>\mathrm{m^{2}\cdot Hz^{-1}},$ (46)
@@
Arm $\text{Arm length}~{}L~{}(10^{9}\mathrm{m})$ 2.5 3 0.17 0.05
@@
Integration time $T_{\mathrm{obs}}~{}(\mathrm{year})$ $4.5$ $5$ $5/2$ $4$
Frequency band $(\mathrm{Hz})$ $[10^{-4},1]$ $[10^{-4},1]$ $[10^{-4},1]$ $[10^{-1},10]$
Table 1: Parameters of of several planned SGWDs Colpi et al. (2024); Luo et al. (2020, 2016); Crowder and Cornish (2005).
@@
$g_{a\gamma}^{i}(f)=\sqrt{\frac{S_{i}(f)\,\pi^{2}\nu_{0}^{2}}{\rho_{\mathrm{DM}}\,C_{i}^{2}(f)\,T_{\mathrm{obs}}}}$ (47)
@@
$\mathrm{SNR}=\frac{S_{i}}{N_{i}}.$ (48)
We assume an SNR of 1 to provide a conservative estimate of the interferometer’s performance. It is noteworthy that the ratio of noise PSD between the first- and second-generation TDI combinations is the same as the ratio of the squares of their corresponding factor $C_{i}(f)$. Therefore, the second-generation TDI combinations ($X_{2},\,\alpha_{2},\zeta_{2}$) yield sensitivity estimates that are identical to those of their first-generation counterparts. In Fig. 2, we illustrate the sensitivity estimates of LISA, Taiji, TianQin, and BBO to the ALDM-photon coupling $g_{a\gamma}$ for the first-generation TDI combinations. These estimates are compared with existing constraints derived from CAST Anastassopoulos et al. (2017), SN 1987A Payez et al. (2015), the M87 galaxy Marsh et al. (2017) and the quasar H1821+643 Reynés et al. (2021). The parameters of interest for the sensitivity curves of various SGWDs are summarized in Table. 1.
@@
Figure 2: Sensitivity estimates of various SGWDs to the ALDM-photon coupling $g_{a\gamma}$ for the first-generation TDI conbinations. The constraints from CAST Anastassopoulos et al. (2017), SN 1987A Payez et al. (2015), the M87 galaxy Marsh et al. (2017) and the quasar H1821+643 Reynés et al. (2021) are also shown for comparison.
The comparisons to existing astrophysical and laboratory constraints reveal that the SGWD-based approach holds potential to improve sensitivities to the ALDM-photon coupling strength, particularly in the mass range of $10^{-19}\,\mathrm{eV}$ to $10^{-14}\,\mathrm{eV}$. Notably, the SGWD senstitivity curves dip below the astrophysical constraints in certain mass windows, reaching $g_{a\gamma}\sim 10^{-13}\>\mathrm{GeV^{-1}}$, signify their competitive reach in the search for ALDM.
@@
V Discussion
The proposed OB modification offers a promising route to probe the ALDM-photon coupling without altering the polarization state of inter-spacecraft links. However, the practical implementation of this modification entails a series of intricate engineering challenges that demand meticulous consideration. For instance, the stabilization of additional optical elements is critical to avoid introducing extra phase noise. Specifically, the 50/50 beam splitter ratio for the ALDM interferometer must be carefully maintained to prevent noise arising from power fluctuations, while additional baffling for the auxiliary optical path is required to mitigate stray light effects.
The projected sensitivity curves, presented in the preceding section, provide preliminary assessment of the instrument’s performance, demonstrating the potential of SGWDs to probe ALDM. Despite these promising sensitivity projections, the impact of the modification on the overall OB design cannot be overlooked. The integration of an auxiliary interferometer into the OB introduces challenges not only in maintaining mechanical and thermal stability but also in managing the total light power budget. Systematic errors arising from the auxiliary path, such as those induced by imperfect beam splitting or misalignment, must be carefully quantified and minimized. The p-polarized light induced by the ALDM may be mitigated by the the existing optics of the OB like transimitting beam clipping, requiring dedicated design for the actual light path. Future research should focus on detailed simulations and laboratory demonstrations to validate the preliminary sensitivity estimates and to develop practical methods for mitigating the additional noise contributions introduced by the modification.
@@ VI Conclusion
Acknowledgements.
QGH is supported by the National Key Research and Development Program of China Grant No.2020YFC2201502, grants from NSFC (grant No. 12475065, 12250010). RMY is supported by the National Natural Science Foundation of China under Grant No. 12447131. BXJ is supported by the National Natural Science Foundation of China under Grant No. 12447105.

### v2

@@ 4.3 Expected sensitivity
where $P_{\mathrm{oms}}$ and $P_{\mathrm{acc}}$ are the noise amplitude parameters. For the BBO, the noises PSDs are given by [Corbin:2005ny]
$S_{\mathrm{oms}}^{\mathrm{BBO}}(f)$ $=\frac{2.0\times 10^{-34}}{3L^{2}}\>\mathrm{m^{2}\cdot Hz^{-1}},$ (4.22)
@@
$\text{Arm length}~L~(10^{9}\mathrm{m})$ 2.5 3 0.17 0.05
@@
Integration time $T_{\mathrm{obs}}~(\mathrm{year})$ $4.5$ 55 $5/2$ 44
Frequency band $(\mathrm{Hz})$ $10^{-4},1] [$$10^{-4},1] [$$10^{-4},1] [$$10^{-1},10]$
Table 1: Parameters of of several planned SGWDs [LISA:2024hlh, Luo:2019zal, TianQin:2015yph, Crowder:2005nr].
@@
$g_{a\gamma}^{i}(f)=\sqrt{\frac{S_{i}(f)\,\pi^{2}\nu_{0}^{2}}{\rho_{\mathrm{DM}}\,C_{i}^{2}(f)\,T_{\mathrm{obs}}}}$ (4.23)
@@
$\mathrm{SNR}=\frac{S_{i}}{N_{i}}.$ (4.24)
We assume an SNR of 1 to provide a conservative estimate of the interferometer’s performance. It is noteworthy that the ratio of noise PSD between the first- and second-generation TDI combinations is the same as the ratio of the squares of their corresponding factors $C_{i}(f)$. Therefore, the second-generation TDI combinations ($X_{2},\,\alpha_{2},\zeta_{2}$) yield sensitivity estimates that are identical to those of their first-generation counterparts. In Fig. 2, we illustrate the sensitivity estimates of LISA, Taiji, TianQin, and BBO to the ALDM-photon coupling $g_{a\gamma}$ for the first-generation TDI combinations. These estimates are compared with existing constraints derived from CAST [CAST:2017uph], SN 1987A [Payez:2014xsa], the M87 galaxy [Marsh:2017yvc] and the quasar H1821+643 [Reynes:2021bpe]. The parameters of interest for the sensitivity curves of various SGWDs are summarized in Table. 1.
@@
Figure 2: Sensitivity estimates of various SGWDs to the ALDM-photon coupling $g_{a\gamma}$ for the first-generation TDI combinations. The constraints from CAST [CAST:2017uph], SN 1987A [Payez:2014xsa], the M87 galaxy [Marsh:2017yvc] and the quasar H1821+643 [Reynes:2021bpe] are also shown for comparison.
The comparisons to existing astrophysical and laboratory constraints reveal that the SGWD-based approach holds potential to improve sensitivities to the ALDM-photon coupling strength, particularly in the mass range of $10^{-19}\,\mathrm{eV}$ to $10^{-14}\,\mathrm{eV}$. Notably, the SGWD sensitivity curves dip below the astrophysical constraints in certain mass windows, reaching $g_{a\gamma}\sim 10^{-13}\>\mathrm{GeV^{-1}}$, signifying their competitive reach in the search for ALDM.
@@
5 Discussion
The proposed OB modification offers a promising route to probe the ALDM-photon coupling without altering the polarization state of inter-spacecraft links. However, the practical implementation of this modification entails a series of intricate engineering challenges that demand meticulous consideration. For instance, the stabilization of additional optical elements is critical to avoid introducing extra phase noise. Specifically, the 50/50 beam splitter ratio for the AUXI must be carefully maintained to prevent noise arising from power fluctuations, while additional baffling for the auxiliary optical path is required to mitigate stray light effects.
The projected sensitivity curves, presented in the preceding section, provide a preliminary assessment of the instrument’s performance, demonstrating the potential of SGWDs to probe ALDM. Despite these promising sensitivity projections, the impact of the modification on the overall OB design cannot be overlooked. The integration of an auxiliary interferometer into the OB introduces challenges not only in maintaining mechanical and thermal stability but also in managing the total light power budget. Systematic errors arising from the auxiliary path, such as those induced by imperfect beam splitting or misalignment, must be carefully quantified and minimized. The p-polarized light induced by the ALDM may be mitigated by the existing optics of the OB like transmitting beam clipping, requiring dedicated design for the actual light path. Future research should focus on detailed simulations and laboratory demonstrations to validate the preliminary sensitivity estimates and to develop practical methods for mitigating the additional noise contributions introduced by the modification.
@@ 6 Conclusion
Acknowledgments
The authors would like to thank the anonymous reviewers for their insightful comments and constructive suggestions, which helped to improve the quality of this paper. QGH is supported by the National Key Research and Development Program of China Grant No.2020YFC2201502, grants from NSFC (grant No. 12475065, 12250010). RMY is supported by the National Natural Science Foundation of China under Grant No. 12447131. BXJ is supported by the National Natural Science Foundation of China under Grant No. 12447105.